### v2

@@ 4 A priori estimates and stability for BSDEs
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}}\qquad 0\leq s\leq T,$
Since $g(s,0,\tilde{z})=0$ by construction, the numerator is simply $g(s,Z_{s},\tilde{Z}_{s})$. Moreover, Assumption $(\rm H_{q})$} yields
$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq\int_{\tilde{\Omega}}\left|L_{s}\right|\,d\tilde{\mathbb{P}}\leq K_{z}\big(1+\left|Z_{s}\right|\big),\qquad 0\leq s\leq T.$
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{t\in[0,T]}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$. By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. Girsanov’s theorem then implies that the process
@@
is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, thus we obtain
$Y_{t}=\xi-\int_{t}^{T}Z_{s}\!\cdot dW_{s}^{L}.$
Hence,
$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T.$
Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely by Lemma 4.3, and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
Then we have a direct result from the previous two propositions.
@@
Let $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.6), if either
@@
We only sketch how the two estimates interact, as the details follow directly from Propositions 4.2 and 4.4. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
@@
For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$ solve
@@
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. Then there exist $p^{*}>1$ and, for every $p>p^{*}$, a constant $C_{p}>0$ such that
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2p}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{1}-\xi^{2}\|_{L^{2p}},$
where $p^{*}>1$ and $C_{p}$ depend only on $T$, $K_{z}$, $K_{\tilde{z}}$, $K_{0}$, and do not depend on $\|Z^{i}\|_{\mathbb{S}^{\infty}}$.
@@
Then we obtain
$\Delta Y_{t}=\Delta\xi+\int_{t}^{T}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}\Delta Z_{s}\cdot dW_{s},\qquad t\in[0,T].$ (4.17)
Preliminary step (linearization, RHI and sliceability). As in the proof of Proposition 4.4, we use a linearization argument: there exist processes $L_{s}$ and $\tilde{L}_{s}$ on the product space such that, for all $s\in[0,T]$,
$L_{s}:=\frac{g(s,Z^{1}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}},\qquad\tilde{L}_{s}:=\frac{g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{2}_{s})}{|\Delta\tilde{Z}_{s}|^{2}}\,\Delta\tilde{Z}_{s}\,\mathbf{1}_{\{\Delta\tilde{Z}_{s}\neq 0\}}$
with estimates
$|L_{s}|\leq K_{z}\big(1+|Z_{s}^{1}|+|Z_{s}^{2}|\big),\qquad|\tilde{L}_{s}|\leq K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big),$ (4.18)
and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$.
According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale by (4.18) and symmetry. Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C.$ (4.19)
@@
Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$. In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending only on $C$. Under $\mathbb{P}^{L}$ we can rewrite (4.17) as
$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in[0,T].$ (4.20)
The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies a reverse Hölder inequality by Lemma 2.1: there exists $q^{*}>1$ such that, for all $1\leq q\leq q^{*}$,
@@
$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,$
@@
$q^{*}$ $:=\phi\!\left(\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right),\qquad\phi(x):=\left(1+\frac{1}{x^{2}}\log\frac{2x-1}{2x-2}\right)^{\!1/2}-1,$
@@
$C_{q}^{*}$ $:=2\left(1-\frac{2q-1}{2q-2}\,\exp\!\left\{q^{2}\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+2\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right\}\right)^{-1}\!.$
By boundedness of $\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$, the constants $q^{*},C_{q}^{*}$ depend only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Finally, set
$M_{t}:=\int_{0}^{t}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s},\qquad t\in[0,T].$
Then $M\in\mathcal{H}^{\infty}(\mathbb{P})$, and by schachermayer1996institut it is sliceable. Furthermore, since $Z^{1},Z^{2}\in\mathbb{S}^{\infty}$, we can (by (jackson2022triangle, Remark 2.8)) construct a deterministic partition $0=t_{0}<t_{1}<\dots<t_{N}=T$ with arbitrarily small mesh size bounded by a prescribed $\varepsilon>0$.
$\mathbb{S}^{p}$–control of $\Delta Y$. Fix $\varepsilon$, $k\in\{0,\dots,N-1\}$ and write $I_{k}=[t_{k},t_{k+1}]$. By subtracting (4.20) at $u=t_{k+1}$ from that at $u\in I_{k}$, we obtain the local BSDE
$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I_{k}.$
Using Cauchy-Schwarz on the copy space, the growth bound (4.18), and then Fefferman’s inequality (Theorem 2.5 in kazamaki2006BMO) on $I_{k}$, we obtain
$\int_{u}^{t_{k+1}}\tilde{\mathbb{E}}|\tilde{L}_{s}||\Delta\tilde{Z}_{s}|ds$ $\leq\mathbb{E}\int_{0}^{T}(\tilde{\mathbb{E}}\big|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)\big|^{2})^{1/2}(\tilde{\mathbb{E}}|\Delta\tilde{Z}_{s}|^{2})^{1/2}ds$ (4.21)

### v3

@@ 4 A priori estimates and stability for BSDEs
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|Z_{s}|^{2}}\,Z_{s}\,\mathbf{1}_{\{Z_{s}\neq 0\}},\qquad 0\leq s\leq T.$
[…]
$|g(s,z,\tilde{z})-g(s,0,\tilde{z})|\leq K_{z}\big(1+|z|+|\tilde{z}|\big)\,|z|.$
Therefore,
$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq K_{z}\big(1+|Z_{s}|+{\mathbb{E}}|Z_{s}|\big),\qquad 0\leq s\leq T.$ (4.18)
Define the process
$M_{t}:=\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s},\qquad t\in[0,T].$
We claim that $M$ is a BMO martingale. Indeed, by (4.18) and $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$,
$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|^{2}\leq 3K_{z}^{2}\Big(1+|Z_{s}|^{2}+{\mathbb{E}}|Z_{s}|^{2}\Big).$
Hence, for any stopping time $\tau$,
${\mathbb{E}}\!\Big[\langle M\rangle_{T}-\langle M\rangle_{\tau}\,\big|\,\mathcal{F}_{\tau}\Big]$ $={\mathbb{E}}\!\Big[\int_{\tau}^{T}\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$
$\leq 3K_{z}^{2}\Big(T+{\mathbb{E}}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]+\int_{0}^{T}{\mathbb{E}}|Z_{s}|^{2}ds\Big).$
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$, and $\mathbb{H}^{2}\subset\mathbb{H}^{2}_{\mathrm{BMO}}$. Therefore $\|M\|_{\mathrm{BMO}}<\infty$, proving the claim. By Kazamaki’s condition, the stochastic exponential $\mathcal{E}(M)$ is a uniformly integrable martingale. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}(M)_{T}\,d\mathbb{P}$; then $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. By Girsanov’s theorem, the process
@@
is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, and (4.17) becomes
$Y_{t}=\xi-\int_{t}^{T}Z_{s}\cdot dW_{s}^{L},\qquad 0\leq t\leq T.$
Thus
$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T,$
and consequently $|Y_{t}|\leq\|\xi\|_{L^{\infty}}$ for all $t\in[0,T]$, $\mathbb{P}^{L}$–a.s. Since $\mathbb{P}^{L}\sim\mathbb{P}$, the same bound holds $\mathbb{P}$–a.s., and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
We now state a direct consequence of the previous two propositions.
@@
Let $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.6), if either
@@
We only sketch how the two estimates interact here. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
@@
Corollary 4.5.
Assume $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is a solution to the BSDE (2.6) with $\xi$ satisfying:
  • $({\rm H_{\xi}})$
    There is a process $H\in\mathbb{H}^{\infty}$ satisfying
    $\xi=\mathbb{E}\xi+\int_{0}^{T}H_{s}dW_{s}.$
Then $Z\cdot W$ is sliceable and moreover, $\left\|\int_{0}^{T}|Z_{s}|^{2}ds\right\|_{L^{\infty}}<\infty.$
Proof.
From the proof of Proposition 4.3, we have
$Y_{t}=\xi-\int_{t}^{T}Z_{s}\cdot dW^{L}_{s}.$
In other words,
$\int_{0}^{t}Z_{s}\cdot dW^{L}_{s}=Y_{t}-Y_{0}=\mathbb{E}^{L}_{t}[\xi]-\mathbb{E}^{L}[\xi].$
Since the quadratic variation does not change under an equivalent change of measure, we have
$\left\|\int_{0}^{T}|Z_{s}|^{2}ds\right\|_{L^{\infty}}=\left\|\int_{0}^{T}|H_{s}|^{2}ds\right\|_{L^{\infty}}<\infty.$
Since the class of sliceable BMO martingales is the BMO closure of martingales with bounded quadratic variation, this shows that $Z\cdot W$ is sliceable. ∎
@@
For $i=1,2$, consider $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
@@
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0}),(\mathrm{H}_{q})$ and $(\mathrm{H}_{\xi})$. Then there exists a constant $C>0$ such that
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{\infty}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}\ \leq\ C\;\|\xi^{1}-\xi^{2}\|_{L^{\infty}},$
where $C$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$ and the quadratic variation bounds of $\mathbb{E}_{t}[\xi^{1}]$ and $\mathbb{E}_{t}[\xi^{2}]$ in Corollary 4.5.
@@
We consider a sequence of stopping times $0=\tau_{0}\leq\tau_{1}\leq\cdots\leq\tau_{N}=T$ that slices the BMO martingales $Z^{1}\cdot W$ and $Z^{2}\cdot W$ into pieces with BMO norm of size $\varepsilon$ and such that $\tau_{j+1}-\tau_{j}<\varepsilon$ for all $j=0,1,\dots,N-1$. Then, for $\tau_{j}\leq t\leq\tau_{j+1}$, we obtain
$\Delta Y_{t}=\Delta Y_{\tau_{j+1}}+\int_{t}^{\tau_{j+1}}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{\tau_{j+1}}\Delta Z_{s}\cdot dW_{s}.$ (4.20)
We define the sliced BMO norm on $\tau_{j},\tau_{j+1}]$ by
$\big\|Z^{i}\cdot W\big\|_{\mathrm{BMO}_{j}(\mathbb{P})}:=\sup_{\tau\in\mathcal{T}}\Big\|\mathbb{E}_{\tau}\Big[\int_{\tau}^{\tau_{j+1}}|Z^{i}_{s}|^{2}\,ds\Big]\Big\|_{L^{\infty}}^{1/2},$
@@
where $\mathcal{T}$ denotes the collection of all stopping times valued in $\tau_{j},\tau_{j+1}]$. Similarly, $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$, we set
$\big\|Z^{i}\cdot W^{L}\big\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}:=\sup_{\tau\in\mathcal{T}_{j}}\Big\|\mathbb{E}_{\tau}^{\mathbb{P}^{L}}\Big[\int_{\tau}^{\tau_{j+1}}|Z^{i}_{s}|^{2}\,ds\Big]\Big\|_{L^{\infty}}^{1/2}.$
According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale. Moreover, by Corollary 4.4 and Lemma 2.2, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$ $+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C,$ (4.21)
@@
$\frac{1}{C}\left\|Z^{i}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$ $\leq\left\|Z^{i}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}\leq C\left\|Z^{i}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})},\qquad i=1,2,\ j=0,1,\dots,N-1,$
Under $\mathbb{P}^{L}$ we can rewrite (4.20) as
$\Delta Y_{t}=\Delta Y_{\tau_{j+1}}+\int_{t}^{\tau_{j+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{t}^{\tau_{j+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad\tau_{j}\leq t\leq\tau_{j+1}.$ (4.22)
By the Itô formula, for any $t\in[\tau_{j},\tau_{j+1}]$, we have
$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+2\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Y_{s}|\left|\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\right|ds$ (4.23)
$\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+2\left\|\sup_{s\in[t,\tau_{j+1}]}|\Delta Y_{s}|\right\|_{L^{\infty}}\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]ds.$
Note that there exists a constant $A$ such that
$\left|\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\right|$ $\leq A\widetilde{\mathbb{E}}[(1+|Z^{1}_{s}|+|Z^{2}_{s}|+|\tilde{Z}^{1}_{s}|+|\tilde{Z}^{2}_{s}|)|\Delta\tilde{Z}_{s}|]$
$\leq A\left[\left(|Z^{1}_{s}|+|Z^{2}_{s}|\right)\mathbb{E}|\Delta Z_{s}|+\mathbb{E}[(1+|Z^{1}_{s}|+|Z^{2}_{s}|)|\Delta Z_{s}|]\right].$
The Fefferman’s inequality for BMO martingales yields that
$\int_{t}^{\tau_{j+1}}\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$ $\leq\sqrt{2}\mathbb{E}\left[\left(\int_{t}^{\tau_{j+1}}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}\right]\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}$
$\leq\sqrt{6}C\left(\varepsilon+\left\|Z^{1}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}+\left\|Z^{2}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}\right)\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
$\leq 3\sqrt{6}C\varepsilon\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}.$
On the other hand,
$\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}\left(|Z^{1}_{s}|+|Z^{2}_{s}|\right)\mathbb{E}|\Delta Z_{s}|ds$ $\leq\sqrt{2}\mathbb{E}^{L}_{t}\left(\int_{t}^{\tau_{j+1}}\left(|Z^{1}_{s}|^{2}+|Z^{2}_{s}|^{2}\right)ds\right)^{1/2}\left(\int_{t}^{\tau_{j+1}}\mathbb{E}|\Delta Z_{s}|^{2}ds\right)^{1/2}$
$\leq\sqrt{2}C\left(\left\|Z^{1}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}+\left\|Z^{2}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}\right)\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
$\leq 2\sqrt{2}C^{2}\varepsilon\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}.$
As a result, (4.23) becomes
$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+4A(3\sqrt{6}C+2\sqrt{2}C^{2})\varepsilon\left\|\sup_{s\in[t,{\tau_{j+1}}]}|\Delta Y_{s}|\right\|_{L^{\infty}}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
$\leq\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left(\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}\right),$
by letting $\varepsilon$ small enough. Then, we have $\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}$ and $\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}$ are both bounded by
$\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}.$
This implies that
$\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}\leq 4\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}.$
Applying the above estimates iteratively from $T$ to 0, we obtain the estimate
$\left\|\sup_{s\in[0,T]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}(\mathbb{P}^{L})}^{2}\leq N4^{N}\left\|\Delta\xi\right\|_{L^{\infty}}^{2},$
where $N$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Another application of Lemma 2.2 proves the claim. ∎
We can remove $\mathrm{(H_{\xi})}$ under a more restrictive structural assumption on $g$.
Theorem 4.7.
For $i=1,2$, consider $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$. Then there exists a constant $C>0$ such that
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{\infty}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}\ \leq\ C\;\|\xi^{1}-\xi^{2}\|_{L^{\infty}},$
where $C$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$.
Proof.
In the previous proof, we can take $\tau_{j}=\frac{jT}{N}$, where $N$ is large enough. Consider $u=\tau_{j}$ and $v=\tau_{j+1}$. By the Itô formula, for any $t\in[u,v]$, we have
$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Y_{s}|\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$
$\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\mathbb{E}^{L}_{t}\sup_{s\in[t,v]}|\Delta Y_{s}|\int_{t}^{v}\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$
$\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\left(\mathbb{E}^{L}_{t}\sup_{s\in[u,v]}|\Delta Y_{s}|\right)\mathbb{E}\int_{u}^{v}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds.$ (4.24)
On the other hand, by the Fefferman’s inequality for BMO martingale, we have
$\mathbb{E}\int_{u}^{v}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds\leq\sqrt{2}\mathbb{E}\left[\left(\int_{u}^{v}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}\right]\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})},$
@@
$\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})}=\left(\sup_{\tau\in\mathcal{T}}\left\|\mathbb{E}_{\tau}\int_{\tau}^{v}|\Delta Z|^{2}ds\right\|_{L^{\infty}}\right)^{1/2}$
and $\mathcal{T}$ denotes the collection of all stopping times valued in $[u,v]$. By Lemma 2.2 applied on the interval $[u,v]$ and Corollary 4.4, there exists a constant $C$, which only depends on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that